# Balancing Accuracy and Efficiency for Large-Scale SLAM: A Minimal Subset Approach for Scalable Loop Closures

### v2

@@ IV-A Quantitative and Qualitative Analysis
IV-B Larger-Scale Evaluations
@@
Metrics t. ATE R. ATE FPR MEM. TIME
Apollo-SB Sunnyvale All Samples nan nan nan >128GB nan
Constant 1.0m 98.5% 98.4% 12% 90.9GB 7340s
Constant 2.0m 98.1% 97.3% 17% 47.2GB 2264s
Constant 3.0m 97.9% 97.0% 28% 35.4GB 1515s
Entropy-based 95.5% 93.4% 15% 51.0GB 4945s
Spaciousness 94.9% 97.5% 24% 38.3GB 1752s
MSA (default) 98.3% 94.1% 12% 41.2GB 1951s
MSA (w/o $\pi_{\tau}$) 94.7% 93.2% 13% 32.4GB 1438s
MSA (w/o $\rho_{\tau}$) 98.9% 96.4% 12% 43.7GB 2022s
MSA ($\alpha=10$) 96.5% 93.9% 12% 37.6GB 1764s
MSA ($\beta=10$) 98.6% 95.8% 12% 42.0GB 2004s
Newer-College Park All Samples 93.9% 78.9% 02% 65.8GB 6742s
Constant 0.1m 89.7% 71.4% 03% 63.2GB 4987s
Constant 0.3m 79.9% 58.5% 08% 28.9GB 1262s
Constant 1.0m 60.7% 43.1% 23% 09.2GB 165s
Entropy-based 87.3% 65.8% 05% 26.3GB 1401s
Spaciousness 72.5% 52.5% 14% 27.2GB 1096s
MSA (default) 90.1% 67.0% 02% 38.7GB 1884s
MSA (w/o $\pi_{\tau}$) 81.7% 41.4% 03% 12.8GB 346s
MSA (w/o $\rho_{\tau}$) 97.7% 81.4% 02% 64.7GB 6226s
MSA ($\alpha=10$) 86.9% 64.4% 02% 29.2GB 1498s
MSA ($\beta=10$) 93.2% 69.4% 02% 42.5GB 2791s
To support our last contribution, we conduct more realistic, larger-scale, online experiments. Poses are processed sequentially, simulating real-time operation. Keyframes are sampled online using the previously mentioned methods, appended to the pose graph, and optimized at each step with iSAM2 [8]. Simultaneously, OT descriptors are queried against past keyframes to detect loop closures, which are verified with small_gicp and integrated into the graph.
IV-B1 Urban Driving Scenario
For this scenario we use one of the datasets from the batch experiments, Apollo-SouthBay, focusing on the SunnyvaleBig- loop sequence, which spans over 100 kilometers. Table II reports the translation and rotation ATE, FPR, total allocated memory, and processing time. Notably, when using all samples, the system exhausts memory, approximately 70% into the sequence, preventing the calculation of the metrics. The other sampling methods reduce memory usage while maintaining excellent performance. This is likely due to the sequence revisiting the same locations multiple times, enabling frequent loop closure edges that mitigate accumulated drift. In Fig. 5, the optimized trajectory of the entropy-based method is compared to MSA, with a color gradient representing translational RPE. The visualization highlights MSA’s superior performance, achieved with lower memory allocation and a reduced false positive rate. Lastly, Fig. 6 shows the memory usage over time, illustrating how sampling methods scale more efficiently compared to retaining all samples or using a constant 1-meter interval.
@@ IV-B1 Urban Driving Scenario
Figure 5: Trajectory comparisons. The ground truth and raw KISS-ICP poses compared against the sampled and optimized poses using the proposed Minimal Subset Approach (MSA) and the entropy-based method for the Sunnyvale and the constant 1-meter approach for the Park. The color gradient represents the Relative Pose Error (RPE). Grid spacing corresponds to 0.5x0.5 km2 for the left figure and 50x50 m2 for the right figure.
IV-B2 Campus Scenario
To provide additional insights, we evaluate a different dataset, the Newer-College [36], and more specifically the Park sequence, a 30-minute mapping session with multiple loops. This scenario, resembling a mobile ground robot, differs significantly from Apollo-SouthBay due to its more confined environment and denser frame sampling, approximately 0.1 meters compared to 1 meter. These differences highlight how distinct missions impose unique requirements.
From Table II, constant intervals show that, a 1-meter sampling, severely underperforms in this scenario, retaining only 7% of the frames. This leads to significant performance losses in both ATE and FPR. To address this, we include lower constant intervals of 0.1 and 0.3 meters for comparison. While these improve performance, our proposed MSA outperforms them by prioritizing informative sampling, providing more valuable nodes to the pose graph. The entropy-based method closely follows, emphasizing the importance of spatial features in sampling. Here, we note that both the spaciousness- and entropy-based methods required additional tuning for this scenario. As in the previous experiment, Fig. 5 compares the optimized trajectory of MSA with the 1-meter approach, using a color gradient to depict translational RPE, which highlights MSA’s superior metric localization performance. Figs 6 and 7 further compare memory allocation over time and processing times. Fig. 7 presents box plots for incremental PGO (left) and Loop Closure Detection (right) processing times per step. Notably, both tasks exhibit significantly high processing times, potentially causing delays in real-world missions. While lower fixed intervals (0.1 and 0.3 meters) reduce processing time, 0.3-meter and 1-meter intervals show subpar performance, and the 0.1-meter interval demands substantial memory allocation since it is very close to the native sampling of the dataset. Interestingly, methods achieving good ATE performance, such as the entropy-based and MSA, demonstrate higher outliers in the incremental PGO, suggesting that valuable edges are added to the pose graph, contributing to their better performance.
@@ IV-C Ablation Study
In all experiments, we maintained constant hyperparameters ($\alpha=\beta=1$), demonstrating that the proposed method adapts seamlessly to different environments and descriptors without requiring manual tuning. To further analyze the impact of each optimization term, we conducted an ablation study, with results shown in Table II. Specifically, we tested scenarios where only one term, $\rho_{\tau}$ or $\pi_{\tau}$, was used, and scenarios where the weight of one term was increased to 10 while keeping the other at 1. The results indicate that emphasizing the information preservation term or using it exclusively improves performance but leading to increased number of keyframes sampled. Conversely, prioritizing redundancy minimization significantly reduces memory usage but negatively impacts performance. Based on these, practitioners can decide whether to prioritize memory or performance, depending on their specific application requirements.
V Discussion
@@
The results highlight that no universal constant sampling interval is suitable for all scenarios. While a constant 1-meter interval performs well in datasets similar to the Urban Driving Scenario, showing minimal deviations in Fig. 3, it performed poorly in the mobile robotics Campus Scenario. This reinforces our core argument: adaptive methods that account for both place recognition performance and pose graph optimization are essential. Naive methods that rely on spatial heuristics, such as spaciousness- and entropy-based approaches, require extensive tuning and prior environmental knowledge, limiting their applicability. Our proposed MSA consistently excels across all scenarios without the need for parameter tuning, demonstrating robustness in this traditional loop closure setup. It is effective with both widely used handcrafted and learning-based descriptors, showcasing adaptability to different datasets and descriptors. As future work, we aim to test our method with full global localization pipelines, such as RING++ [37] or LCDNet [38].
VI Conclusions
In conclusion, as the robotics community advances towards larger-scale and long-term missions, addressing system limitations with efficient and scalable solutions becomes crucial. This article introduced the Minimal Subset Approach (MSA) to address the combined challenges of loop closure detection and pose graph optimization by reducing redundant samples while maintaining robust performance. Experimental evaluations across multiple datasets show that MSA excels in terms of FPR, as well as in achieving superior ATE and RPE after pose graph optimization. Unlike other adaptive sampling methods, MSA requires minimal to no parameter tuning, while using a fraction of the system’s memory and maintaining comparable computational times.

### v3

@@ V-A Quantitative and Qualitative Analysis
V-B Larger-Scale Evaluations
@@
Metrics t. ATE R. ATE F1-MAX MEM. TIME
Apollo-SB Sunnyvale
All Samples nan nan nan >128GB nan
Constant 1.0m 98.5% 98.4% 93.3% 90.9GB 7340s
Constant 2.0m 98.1% 97.3% 62.1% 47.2GB 2264s
Constant 3.0m 97.9% 97.0% 46.3% 35.4GB 1515s
Entropy-based 95.5% 93.4% 81.9% 51.0GB 4945s
Spaciousness 94.9% 97.5% 46.1% 38.3GB 1752s
MSA (default) 98.2% 97.5% 94.4% 40.3GB 1923s
MSA (w/o $\pi_{\tau}$) 94.8% 93.2% 90.1% 32.4GB 1473s
MSA (w/o $\rho_{\tau}$) 98.9% 97.7% 94.8% 43.9GB 2124s
MSA ($\alpha=10$) 98.6% 97.7% 94.5% 49.2GB 2889s
MSA ($\beta=10$) 96.1% 94.7% 90.5% 34.9GB 1572s
Newer-College Park
All Samples 93.9% 78.9% 99.8% 65.8GB 6742s
Constant 0.1m 89.7% 71.4% 99.8% 63.2GB 4987s
Constant 0.3m 79.9% 58.5% 94.8% 28.9GB 1262s
Constant 1.0m 60.7% 43.1% 64.5% 09.2GB 165s
Entropy-based 87.3% 65.8% 92.6% 26.3GB 1401s
Spaciousness 72.5% 52.5% 89.0% 27.2GB 1096s
MSA (default) 92.9% 73.1% 96.9% 40.7GB 1925s
MSA (w/o $\pi_{\tau}$) 85.9% 69.4% 93.7% 32.2GB 1482s
MSA (w/o $\rho_{\tau}$) 93.2% 73.1% 97.8% 44.0GB 2191s
MSA ($\alpha=10$) 93.7% 75.5% 98.2% 54.8GB 2352s
MSA ($\beta=10$) 88.3% 65.9% 89.5% 35.1GB 1749s
The ATEs are presented as the percent improv. in the trajectory after PGO.
To support our last contribution, that MSA achieves higher localization and place recognition performance than other adaptive methods, while reducing memory usage and computation time relative to the baseline, we conduct more realistic, larger-scale, online experiments. Poses are processed sequentially, simulating real-time operation. Keyframes are sampled online using the previously mentioned methods, appended to the pose graph, and optimized at each step with iSAM2 [Kaess2011isam2]. Simultaneously, OT descriptors are queried against past keyframes to detect loop closures, which are verified with small_gicp and integrated into the graph.
V-B1 Urban Driving Scenario
For this scenario we use one of the datasets from the batch experiments, Apollo-SouthBay, focusing on the SunnyvaleBig- loop sequence, which spans over 100 kilometers. Table II reports the translation and rotation ATE improvements, F1-max score, total allocated memory, and processing time. Notably, when using all samples, the system exhausts memory, approximately 70% into the sequence, preventing the calculation of the metrics. The other sampling methods reduce memory usage while maintaining excellent performance. This is likely due to the sequence revisiting the same locations multiple times, enabling frequent loop closure edges that mitigate accumulated drift. In Fig. 6, the optimized trajectory of the entropy-based method is compared to MSA. The visualization highlights MSA’s better performance, achieved with lower memory allocation and a higher F1-max score. Lastly, Fig. 7 shows the memory usage over time, illustrating how sampling methods scale more efficiently compared to retaining all samples or using a constant 1-meter interval.
@@ V-B1 Urban Driving Scenario
Figure 6: Trajectory comparisons. The ground truth and raw KISS-ICP poses compared against the sampled and optimized poses using the proposed Minimal Subset Approach (MSA) and the entropy-based method for the Sunnyvale and the constant 1-meter approach for the Park. The color gradient represents the translational Relative Pose Error (RPE). Grid spacing corresponds to 0.5x0.5 km2 for the left figure and 50x50 m2 for the right figure.
V-B2 Campus Scenario
To provide additional insights, we evaluate a different dataset, the Newer-College [Ramezani2020newercollege], and more specifically the Park sequence, a 30-minute mapping session with multiple loops. This scenario, resembling a mobile ground robot, differs significantly from Apollo-SouthBay due to its more confined environment and denser frame sampling, approximately 0.1 meters compared to 1 meter. These differences highlight how distinct missions impose unique requirements.
From Table II, constant intervals show that, a 1-meter sampling, severely underperforms in this scenario, retaining only 7% of the frames. This leads to significant performance losses in both ATE and FPR. To address this, we include lower constant intervals of 0.1 and 0.3 meters for comparison. While these improve performance, our proposed MSA outperforms them by prioritizing informative sampling, providing more valuable nodes to the pose graph. The entropy-based method closely follows, emphasizing the importance of spatial features in sampling. Here, we note that both the spaciousness- and entropy-based methods required additional tuning for this scenario. As in the previous experiment, Fig. 6 compares the optimized trajectory of MSA with the 1-meter approach, which highlights MSA’s higher metric localization performance. Figs 7 and 8 further compare memory allocation over time and processing times. Fig. 8 presents box plots for incremental PGO (left) and Loop Closure Detection (right) processing times per step. Notably, both tasks exhibit significantly high processing times, potentially causing delays in real-world missions. While lower fixed intervals (0.1 and 0.3 meters) reduce processing time, 0.3-meter and 1-meter intervals show subpar performance, and the 0.1-meter interval demands substantial memory allocation since it is very close to the native sampling of the dataset. Interestingly, methods achieving good ATE performance, such as the entropy-based and MSA, demonstrate higher outliers in the incremental PGO, suggesting that valuable edges are added to the pose graph, contributing to their better performance.
V-C Ablation Study
@@
In all experiments, we maintained constant hyperparameters ($\alpha=\beta=1$ and $\lambda_{l}=0.1$, $\lambda_{u}=3$), demonstrating that the proposed method adapts seamlessly to different environments and descriptors without requiring manual tuning. To further analyze the impact of each optimization term, we conducted an ablation study, with results shown in Table II. Specifically, we tested scenarios where only one term, $\rho_{\tau}$ or $\pi_{\tau}$, was used, and scenarios where the weight of one term was increased to 10 while keeping the other at 1. The results indicate that emphasizing the information preservation term or using it exclusively improves performance but leading to increased number of keyframes sampled. Conversely, prioritizing redundancy minimization significantly reduces memory usage but negatively impacts performance. Based on these, practitioners can decide whether to prioritize memory or performance, depending on their specific application requirements. Additionally, Fig. 9 reports an ablation of the distance bounds, sweeping $\lambda_{l}\!\in[0,2]$ and $\lambda_{u}\!\in[1,10]$, and evaluating all previously discussed metrics. Accuracy remains largely stable, while memory is more sensitive: larger $\lambda_{u}$ reduces memory on Apollo with minimal loss but degrades accuracy on Newer–College. The latter is a more confined environment where salient features are not widely spaced, which induces rapid viewpoint changes; as $\lambda_{u}$ grows, consecutive frames become too far apart for descriptors to remain smooth, weakening the information-preservation term and reducing the reliability of PGO edges. We therefore adopt $\lambda_{l}=0.1$ to avoid near-duplicates and $\lambda_{u}=3$ to preserve descriptor and odometry locality.
Figure 9: Adaptive bounds ablation. For each dataset, we sweep $\lambda_{l}$, while keeping $\lambda_{u}=5.0$, and $\lambda_{u}$ while keeping $\lambda_{l}=0.1$. We report translational and rotational ATE improvement, the F1–max from the precision and recall curves, and the memory allocation compared to all samples.
VI Discussion and Conclusions
@@
The results highlight that no universal constant sampling interval is suitable for all scenarios. While a constant 1-meter interval performs well in urban driving datasets like Apollo-SouthBay, it performed poorly in the mobile robotic scenario of Newer-College. This reinforces our core argument: adaptive methods that account for both place recognition and pose graph optimization are essential. Naive heuristic approaches based on spaciousness or entropy require extensive tuning and prior environmental knowledge, limiting applicability. MSA consistently performs well across scenarios without parameter tuning, and is effective with both widely used hand-crafted and learning-based descriptors. As future work, we plan to integrate MSA with full global localization pipelines such as LCDNet [cattaneo_lcdnet_2022] and RING++ [Xu2023ring].
In conclusion, as robotics scales to larger and longer missions, efficient and scalable sampling becomes crucial. The Minimal Subset Approach reduces redundant samples while maintaining robust performance. Across multiple datasets, MSA delivers better place recognition and superior ATE/RPE than adaptive methods, and compared to using all samples it reduces memory consumption and computation time.